# The unexpected uses of a bowling pin: anisotropic flow in fixed-target 208Pb+20Ne collisions as a probe of quark-gluon plasma

### v1

@@ Abstract
Over the past decades, collider experiments with atomic nuclei have enabled the exploration of the phase diagram of strong-interaction matter over a wide range of temperature and densities. At ultra-relativistic energies, experiments at the Relativistic Heavy Ion Collider (RHIC) and at the CERN Large Hadron Collider (LHC) have established the formation of quark-gluon plasma (QGP) Harris:2024aov , the hot phase of QCD behaving like a near-ideal fluid of deconfined partons, via the observation of anisotropic flow Ollitrault:2023wjk in combination with the modification of hard probes due to their interaction with the hot environment Apolinario:2022vzg . At these energies, anisotropic flow signals persist down to small systems Nagle:2018nvi ; Shen:2020mgh ; Schenke:2021mxx ; Noronha:2024dtq , including $p$+$p$, $\gamma+$Pb, and $e^{+}$+$e^{-}$ collisions  CMS:2016fnw ; ATLAS:2018ngv ; ALICE:2019zfl ; ATLAS:2021jhn ; Zhao:2022ayk ; Chen:2023njr , although without any visible hard-probe modification, leaving the question of the formation of QGP up to debate ATLAS:2014qaj ; ALICE:2016dei . Consequently, research on small systems has shed light on the apparent hydrodynamization of out-of-equilibrium QCD matter Berges:2020fwq ; Soloviev:2021lhs . In parallel, the Beam Energy Scan (BES) at RHIC has elucidated the equation of state and the transport properties of the QGP towards high net-baryon density PHOBOS:2006dbo ; STAR:2016vqt ; STAR:2017idk , while fixed-target experiments close to the GeV scale have demonstrated the dominance of hadronic degrees of freedom and opened connections with the physics of neutron stars HADES:2019auv ; STAR:2021yiu ; HADES:2022osk . This wealth of experimental measurements in combination with advanced Bayesian inference methods Annala:2019puf ; OmanaKuttan:2022aml ; Giacalone:2023cet ; Paquet:2023rfd ; Shen:2023awv has helped shape a picture of nuclear matter gradually acquiring consistency across experiments and theoretical results.
To add to this picture, exciting prospects lie ahead from fixed-target ion-ion collisions at the LHC. The LHCb detector hosts in particular the System for Measuring Overlap with Gas (SMOG2) LHCb:2021ysy ; LHCb:2022qvj ; Mariani:2022klh , enabling us to perform beam-target collisions at relativistic energy in the centre-of-mass frame ($\sqrt{s_{\rm NN}}=68.5$ GeV for beam nucleons boosted to 2.5 TeV). This configuration is of unique interest. At $\sqrt{s_{\rm NN}}=68.5$ GeV, Pb+O or Pb+Ne collisions yield multiplicities close to those of collider-mode proton-nucleus or oxygen-oxygen collisions at LHC energy, and should exhibit visible collective behavior. At the same time, the beam energy is approximately that explored by the BES at RHIC, implying a significantly baryon-rich system. However, the size of the Pb+O and Pb+Ne droplets is somewhat intermediate between the $d$+Au and Au+Au collisions explored at the BES. Fixed-target collisions at LHC probe, then, a combination of size and baryon density which has not been explored (and will not be explored) by any other experiment.
@@
Characterizations of the matter produced in these experiments have only just begun. The LHCb collaboration has recently published the measurement of $J/\Psi$ and $D_{0}$ production in Pb+Ne collisions LHCb:2022qvj . The resulting ratio between the two cross sections is consistent with the $p$+Ne baseline, even in high-multiplicity events, suggesting that hot-medium interactions may not be effective. Experimental studies of the soft sector and the collective flow have not been performed yet. It is crucial to explore if sizable anisotropic flow is observed in these collisions: in the absence of hard-probe modification, this would effectively open a small system puzzle in a new type of system. Surprisingly, theoretical studies of anisotropic flow in fixed-target mode are missing. To motivate and underscore the potential of the experimental effort, in this Letter we aim to fill this important gap.
@@
In summary, we have unveiled the great opportunities offered by studies of the anisotropic flow in fixed-target collisions at the LHCb detector. Our predictions may be amenable to experimental verification already with the next LHC ion run. The availability of the bowling-pin nucleus 20Ne in SMOG2 enables one to isolate strong effects of the hydrodynamic response. If the dramatic enhancements of $v_{2}$ and $v_{4}$ relative to the Pb+O baseline will be confirmed by the LHCb collaboration, they will hint at the validity of a QGP description. This will pave the way to quantitative characterizations of the matter formed in these experiments via future Bayesian analyses.
The predictions of this manuscript can indeed be improved in several directions. The impact of short-range two-particle correlations (non flow) is washed out in our simulations due to the oversampling of the UrQMD events, which can be relaxed in future studies. In addition, a more detailed analysis of the systematics of the theoretical model will elucidate whether theoretical uncertainties cancel when normalizing Pb+Ne observables with those coming from another collision species Giacalone:2024luz .
@@
We thank the LHCb Collaboration for sharing with us the data shown in Fig. 1. We thank in particular Giacomo Graziani and the participants of the program+workshop “Exploring Nuclear Physics Across Energy Scales 2024” for useful discussions. G.G. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 273811115 – SFB 1225 ISOQUANT, and under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under DOE Contract No. DE-SC0012704 (B.P.S.) and Award No. DE-SC0021969 (C.S.), and within the framework of the Saturated Glue (SURGE) Topical Theory Collaboration. C.S. acknowledges a DOE Office of Science Early Career Award. W.B.Z. is supported by DOE under Contract No. DE-AC02-05CH11231, by NSF under Grant No. OAC-2004571 within the X-SCAPE Collaboration, and within the framework of the SURGE Topical Theory Collaboration. This research was done using resources provided by the Open Science Grid (OSG) Pordes:2007zzb ; Sfiligoi:2009cct , which is supported by the National Science Foundation award #2030508. This work is supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC AdG EXOTIC, grant agreement No. 101018170), by DFG and NSFC through funds provided to the Sino-German CRC 110 “Symmetries and the Emergence of Structure in QCD” (NSFC Grant No. 11621131001, DFG Grant No. TRR110). The work of UGM was supported in part by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2018DM0034). R.R. is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Projektnummer 279384907 – SFB 1245. T.R.R. is supported by the Spanish MCIU (PID2021- 127890NB-I00). The PGCM calculations were performed using HPC resources from GENCI-TGCC (Contracts No. A0130513012 and A0150513012) and CCRT (TOPAZE supercomputer). The NLEFT code development and production calculations utilized the following computational resources: the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) and special GPU time allocated on JURECA-DC; the Oak Ridge Leadership Computing Facility through the INCITE award “Ab-initio nuclear structure and nuclear reactions”; and the TUBITAK ULAKBIM High Performance and Grid Computing Center (TRUBA resources). J.N.H. acknowledges financial support from the US-DOE Nuclear Science Grant No. DE-SC0023861 and within the framework of the Saturated Glue (SURGE) Topical Theory.
@@ II Cumulants of flow fluctuations in the lab frame
Information is lost and we end up with a modified Bessel function, $I_{0}$, which is an even function. The cumulants of the angle-averaged distribution of the $\textbf{v}_{n}$ vector, whose standard notation is $v_{n}\{m\}^{m}$, are then obtained from the equality:
$\ln\langle I_{0}(kv_{n})\bigr{\rangle}=\sum_{m=0}^{\infty}c_{m}k^{m}v_{n}\{m\}^{m}.$ (13)
The integer $m\geq 2$ is an even number, while the coefficients $c_{m}$ have to be matched to the power expansion of the left-hand side, that is:
$\ln\langle I_{0}(kv_{n})\bigr{\rangle}=\ln\biggl(1+\frac{\langle v_{n}^{2}\rangle k^{2}}{4}-\frac{\langle v_{n}^{4}\rangle k^{4}}{64}+\frac{\langle v_{n}^{6}\rangle k^{6}}{576}+\ldots\biggr{missing}).$ (14)
@@
$v_{n}\{4\}^{4}$ $=2\langle v_{n}^{2}\rangle-\langle v_{n}^{4}\rangle,$
@@
and so on, where averages are over events at a given centrality. The cumulants are thus expressed in terms of even moments of the distribution of $v_{n}$, of the form $\langle v_{n}^{2k}\rangle$. These moments represent the quantities that can be measured in the experiments from established multi-particle correlation techniques Ollitrault:2023wjk .
@@
This implies the following. The second-order cumulant $v_{2}\{2\}^{2}$ isolates the variance of the two-dimensional distribution. It tells us how a change in the mean square value of the flow vector magnitude, $\langle v_{n}^{2}\rangle$, is associated with an increase in the variances of its $x$ and $y$ components. Concerning the fourth-order cumulant, conventionally denoted by  [^1] The sign takes into account the negative sign of the $k^{4}$ term in the expansion of Eq. (14).
@@ III Understanding the enhancement of $v_{4}$
In central collisions, one expects the term coupling to elliptic flow to be sub-leading compared to the uncorrelated contribution, $\langle v_{\rm 4L}^{2}\rangle\gg\chi_{4}\langle v_{2}^{4}\rangle$. This justifies a Taylor expansion of the previous expression:
@@
Now, we show in Fig. 5 hydrodynamic results for $\chi_{4,22}$ [panel (a)] and $\langle v_{\rm 4L}\{2\}\equiv\sqrt{\langle v_{\rm 4L}^{2}\rangle}$ [panel (b)]. We see that both these quantities are rather universal and receive only percent-level corrections if one varies the nuclear species/structure. For simplicity, let us then consider that
@@
This equality should hold in central collisions if the departure from unity in the quadrangular flow ratio is due to the nonlinear coupling term to the elliptic flow. Plugging in numbers from the 0-5% class, we obtain the results shown in Fig. 5(c). Agreement between the $v_{4}\{2\}$ ratio and the prediction of Eq. (27) is excellent in central collisions, confirming that it is the nonlinear coupling to elliptic flow that drives the nontrivial Pb+Ne to Pb+O ratio.

### v2

# Anisotropic flow in fixed-target 208Pb+20Ne collisions as a probe of quark-gluon plasma

@@ Abstract
Over the past decades, high-energy collisions of atomic nuclei have provided significant insights into the phase diagram of strong-interaction matter across a wide range of temperatures and densities Busza:2018rrf ; Paquet:2023rfd . These investigations also offer a new way to study nuclear structure, allowing a direct view of many-body correlations of nucleons (e.g., deformations) in the ground states of the colliding nuclei, overcoming some limitations of low-energy nuclear structure experiments STAR:2024wgy .
At ultra-relativistic energies, experiments at the Relativistic Heavy Ion Collider (RHIC) and CERN Large Hadron Collider (LHC) have revealed the formation of quark-gluon plasma (QGP) Harris:2024aov , the hot phase of nuclear matter where quarks and gluons are no longer confined inside protons and neutrons. The formation of QGP has been confirmed through the combined observation of anisotropic flow Shen:2020mgh ; Ollitrault:2023wjk , which characterizes the bulk dynamics of the matter, and of the modifications of the hard probes, energetic particles interacting with the hot plasma Apolinario:2022vzg . These effects have been observed in large systems, like Au+Au and Pb+Pb collisions. In smaller systems, like proton-proton (p+p) and proton-nucleus (p+Pb) collisions, anisotropic flow is observed  PHENIX:2018lia ; ATLAS:2014qaj ; CMS:2016fnw ; ATLAS:2018ngv ; ALICE:2019zfl ; ATLAS:2021jhn ; Zhao:2022ayk ; STAR:2022pfn ; Chen:2023njr albeit in the absence of hard-probe modifications ALICE:2016dei ; CMS:2016xef ; ATLAS:2022iyq , leaving the question of small-sized QGP formation open Nagle:2018nvi ; Schenke:2021mxx ; Noronha:2024dtq .
Fresh insights are expected from the LHC’s new fixed-target program. The LHCb detector hosts the System for Measuring Overlap with Gas (SMOG2) LHCb:2021ysy ; LHCb:2022qvj ; Mariani:2022klh , enabling beam-target collisions at relativistic energy in the center-of-mass frame ($\sqrt{s_{\rm NN}}=68.5$ GeV for beam nucleons boosted to 2.5 TeV). Intermediate-sized systems like lead-neon (Pb+Ne) and lead-oxygen (Pb+O) collisions provide a critical middle ground to understanding the transition from a large to a potentially small QGP. Additionally, due to the lower energies compared to, e.g., symmetric O+O collisions, these systems may present higher baryon density, allowing further tests of the equation of state for baryon-rich nuclear matter Annala:2019puf ; OmanaKuttan:2022aml ; Giacalone:2023cet ; Shen:2023awv in system sizes that have not been explored at the RHIC Beam Energy Scan (BES) or elsewhere, which focused on Au+Au collisions HADES:2019auv ; STAR:2021yiu ; HADES:2022osk .
Most importantly, light-heavy ion collisions help reduce uncertainties related to the initial state geometry. Our understanding of the interaction geometry of proton-nucleus, deuteron-nucleus, and 3He-nucleus collisions suffers from a poor knowledge of the internal spatial structure of nucleons at high energy Schenke:2021mxx . With nuclei like 20Ne and 16O, the initial state geometry will be dominated by the nucleon distributions, and we can rely on precise nuclear structure calculations based on effective field theories of low-energy QCD. This leads to a clearer understanding of the initial conditions of the collisions, which is crucial for interpreting the dynamics of the formed matter. In turn, this also allows new experimental tests of the predicted nuclear shapes Giacalone:2024luz and alpha-clustering YuanyuanWang:2024sgp ; Zhang:2024vkh in the ground states of these isotopes.
@@
Characterizations of the matter produced in fixed-target experiments have only just begun. The LHCb collaboration has recently published the measurement of $J/\Psi$ and $D_{0}$ production in Pb+Ne collisions LHCb:2022qvj . The resulting ratio between the two cross sections is consistent with the $p$+Ne baseline, even in high-multiplicity events, suggesting that hot-medium interactions may not be effective. Therefore, it is crucial to explore if sizable anisotropic flow is observed in these collisions. Experimental studies of the soft sector and the collective flow have not been performed yet, while, surprisingly, theoretical studies are also missing. To motivate and underscore the potential of the experimental effort, in this Letter we aim to fill this important gap.
@@
In summary, we have unveiled the great opportunities offered by studies of the anisotropic flow in fixed-target collisions at the LHCb detector. Our predictions may be amenable to experimental verification already with the next LHC ion run. The availability of the bowling-pin nucleus 20Ne in SMOG2 enables one to isolate strong effects of the hydrodynamic response. If the dramatic enhancements of $v_{2}$ and $v_{4}$ relative to the Pb+O baseline will be confirmed by the LHCb collaboration, they will hint at the validity of a QGP description. This will pave the way to quantitative characterizations of the matter formed in these experiments via future Bayesian analyses in hydrodynamic calculations. Moreover, it would be insightful to compare these hydrodynamic results to those obtained from alternative dynamical approaches, such as transport Lin:2021mdn , based on the same ab initio nuclear input. This will permit us to more precisely pin down the microscopic mechanisms underlying the three-dimensional flow produced in the ion-gas collisions, with good control over the interaction geometry.
The predictions of this manuscript can be improved in several directions. The impact of short-range two-particle correlations (non flow) is washed out in our simulations due to the oversampling of the UrQMD events, which can be relaxed in future studies. In addition, a more detailed analysis of the systematics of the theoretical model will elucidate whether theoretical uncertainties cancel when normalizing Pb+Ne observables with those coming from another collision species Giacalone:2024luz .
@@
We thank the LHCb Collaboration for sharing with us the data shown in Fig. 1. We thank in particular Giacomo Graziani and the participants of the program+workshop “Exploring Nuclear Physics Across Energy Scales 2024” for useful discussions. We thank Mubarak Alqahtani for a careful reading of the manuscript. G.G. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 273811115 – SFB 1225 ISOQUANT, and under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under DOE Contract No. DE-SC0012704 (B.P.S.) and Award No. DE-SC0021969 (C.S.), and within the framework of the Saturated Glue (SURGE) Topical Theory Collaboration. C.S. acknowledges a DOE Office of Science Early Career Award. W.B.Z. is supported by DOE under Contract No. DE-AC02-05CH11231, by NSF under Grant No. OAC-2004571 within the X-SCAPE Collaboration, and within the framework of the SURGE Topical Theory Collaboration. This research was done using resources provided by the Open Science Grid (OSG) Pordes:2007zzb ; Sfiligoi:2009cct , which is supported by the National Science Foundation award #2030508. This work is supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC AdG EXOTIC, grant agreement No. 101018170), by DFG and NSFC through funds provided to the Sino-German CRC 110 “Symmetries and the Emergence of Structure in QCD” (NSFC Grant No. 11621131001, DFG Grant No. TRR110). The work of UGM was supported in part by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2018DM0034). R.R. is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Projektnummer 279384907 – SFB 1245. T.R.R. is supported by the Spanish MCIU (PID2021- 127890NB-I00). The PGCM calculations were performed using HPC resources from GENCI-TGCC (Contracts No. A0130513012 and A0150513012) and CCRT (TOPAZE supercomputer). The NLEFT code development and production calculations utilized the following computational resources: the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) and special GPU time allocated on JURECA-DC; the Oak Ridge Leadership Computing Facility through the INCITE award “Ab-initio nuclear structure and nuclear reactions”; and the TUBITAK ULAKBIM High Performance and Grid Computing Center (TRUBA resources). J.N.H. acknowledges financial support from the US-DOE Nuclear Science Grant No. DE-SC0023861 and within the framework of the Saturated Glue (SURGE) Topical Theory.
@@ II Cumulants of flow fluctuations in the lab frame
Information is lost and we end up with a modified Bessel function, $I_{0}$, which is an even function. The cumulants of the angle-averaged distribution of the $\textbf{v}_{n}$ vector, whose standard notation is $v_{n}\{m\}^{m}$, are obtained from the equality:
$\ln\langle I_{0}(kv_{n})\bigr{\rangle}=\sum_{m=2}^{\infty}c_{m}k^{m}v_{n}\{m\}^{m},$ (13)
where $c_{m}=0$ for odd $m$, while $c_{2}=1/4$, $c_{4}=-1/64$, $c_{6}=1/576$ Bhalerao:2018anl . Expanding the left-hand side of the previous equation,
$\ln\langle I_{0}(kv_{n})\bigr{\rangle}=\ln\biggl(1+\frac{\langle v_{n}^{2}\rangle k^{2}}{4}+\frac{\langle v_{n}^{4}\rangle k^{4}}{64}+\frac{\langle v_{n}^{6}\rangle k^{6}}{2304}+\ldots\biggr{missing}),$ (14)
@@
$v_{n}\{4\}^{4}$ $=2\langle v_{n}^{2}\rangle^{2}-\langle v_{n}^{4}\rangle,$
@@
where averages are over events at a given centrality. The cumulants are thus expressed in terms of even moments of the distribution of $v_{n}$, of the form $\langle v_{n}^{2k}\rangle$. These moments represent the quantities that can be measured in the experiments from established multi-particle correlation techniques Ollitrault:2023wjk .
@@
This implies the following. The second-order cumulant $v_{2}\{2\}^{2}$ isolates the variance of the two-dimensional distribution. It tells us how a change in the mean square value of the flow vector magnitude, $\langle v_{n}^{2}\rangle$, is associated with an increase in the variances of its $x$ and $y$ components. Concerning the fourth-order cumulant, conventionally denoted by  [^1] The sign comes from the negative sign of the $k^{4}$ term in the expansion of Eq. (14).
@@ III Understanding the enhancement of $v_{4}$
In central collisions, one expects the term coupling to elliptic flow to be sub-leading compared to the uncorrelated contribution, $\langle v_{\rm 4L}^{2}\rangle\gg\chi_{4}^{2}\langle v_{2}^{4}\rangle$. This justifies a Taylor expansion of the previous expression:
@@
Now, we show in Fig. 5 hydrodynamic results for $\chi_{4}$ [panel (a)] and $v_{\rm 4L}\{2\}\equiv\sqrt{\langle v_{\rm 4L}^{2}\rangle}$ [panel (b)]. We see that both these quantities are rather universal and receive only percent-level corrections if one varies the nuclear species/structure. For simplicity, let us then consider that
@@
This equality holds in central collisions if the departure from unity in the quadrangular flow ratio is due to the nonlinear coupling term to the elliptic flow. Plugging in numbers from the 0-5% class, we obtain the results shown in Fig. 5(c). Agreement between the $v_{4}\{2\}$ ratio and the prediction of Eq. (27) is excellent in central collisions, confirming that it is the nonlinear coupling to elliptic flow that drives the nontrivial Pb+Ne to Pb+O ratio.